Mack Industries' free cash flow last year was $1 million (i.e., FCF0 = $1 million). You project the company's free cash flow to grow 20% this year (i.e., FCF1 = $1.2 million) and 15% next year. After two years its free cash flow is expected to grow at a constant rate of 5%. The cost of capital is 12%. What is the company's present value of operations (VOP)?

A. $12.23 million

B. $18.67 million

C. $16.65 million

D. $16.91 million
Answer #1

FCF1=$1.2 million

FCF2=(1.2*1.15)=$1.38million

Value after year 2=(FCF2*Growth Rate)/(Cost of capital-Growth Rate)

=(1.38*1.05)/(0.12-0.05)

=$20.7million

Hence VOP=Future FCF*Present value of discounting factor(12%,time period)

=1.2/1.12+1.38/1.12^2+20.7/1.12^2

which is equal to

=$18.67 million(Approx).